OpenStudy (spring98):

@neonmeanings Variable p is 2 more than variable d. Variable p is also 1 less than variable d. Which pair of equations best models the relationship between p and d? A. p = d + 2 p = d – 1 B. p = d – 2 p = d + 1 C. d = 2p d = 2p – 1 D. d = 2p d = 2p + 1

OpenStudy (spring98):

8 th

OpenStudy (anonymous):

you should ask @perl he or she gives help

OpenStudy (spring98):

do you go to CVA?

OpenStudy (anonymous):

no my sister goes to that one i go to MGLVA

OpenStudy (spring98):

ohh nice

OpenStudy (spring98):

@perl can you help me?

OpenStudy (anonymous):

Ok well i hope they help u

OpenStudy (perl):

sure, we can translate the english into algebra

OpenStudy (perl):

Variable p is 2 more than variable d. `Variable p` means p `is` means equals "=" `2 more than` means add 2 or plus 2 p = 2 + d

OpenStudy (perl):

does that make sense?

OpenStudy (spring98):

yes thanks alot! you actually do explain good!

OpenStudy (perl):

for the second part "Variable p is 1 less than variable d." `Variable p is` means "p =" We want one less than variable d, that means `d -1` Combining this we have: p = d - 1